Circular No. 5739

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 R. M. Humphreys, G. S. Aldering, C. O. Bryja and P. M. Thurmes,

University of Minnesota, communicate: “We have a probable identification of the progenitor star from scans of U, B, V and R Kitt Peak 4-m prime-focus plates with the Automated Plate Scanner at the University of Minnesota. These plates were originally obtained on 1982 Mar. 21-23 by Humphreys and were scanned for a stellar content study of M81 by Zickgraf and Humphreys (1991, A.J. 102, 113). Inspecting these scans and using the astrometric solution from the above survey, we find a mag 21 star at α = 9^{h}51^{m}19^{s}.2, δ = +69^{o}15′26′′.5 (equinox 1950.0), in agreement with the supernova position on IAUC 5731. As additional confirmation of this identification, we used an I-band CCD image of SN 1993J, made available by M. Richmond over internet, to measure the positions of six reference stars in the immediate vicinity of SN 1993J using imexamine in IRAF. The corresponding positions on the 4-m plates were determined from the isodensity scans made with the APS. A linear transformation between these sets of coordinates was then calculated. The root-mean-square uncertainty of the transformation was < 0′′.15 in each coordinate. The candidate progenitor has a position on all four 4-m plates that is within 0′′.15 of the position of SN 1993J on the CCD frame. The magnitudes measured from these plates are: V = 20.8±0.1, B = 21.4±0.2, R = 19.9±0.2 and U = 21.7±0.2. Our V is 0.8 mag fainter than reported by Perelmuter on IAUC 5736. Thus the star may be variable; all of our photometry is from plates obtained within two days of each other. The colors are consistent with the candidate progenitor being a highly reddened late A or early F-type supergiant. We have looked for evidence of blending in our images, but as well as can be determined from our plates (1′′-2′′ seeing) the image is isolated. The star is in an association of faint blue stars as seen on our U and B plates. It is also on the inside edge of a prominent dust lane or cloud, so reddening should be considered when interpreting the photometry.”

M. Richmond, Princeton University, reports: “The progenitor of SN 1993J could not have been a blue supergiant. Images of M81 obtained in the ultraviolet (Hill et al. 1992, Ap.J. 395, L37) and in Hα (Petit et al. 1988, A.Ap.Suppl. 74, 475) show no H II region within ∼ 10′′ (= 150 pc) of the supernova’s position. The Hα result puts an upper limit on the luminosity of any H II region in the area that would be exceeded by a main sequence O star. The ultraviolet images also indicate that the extinction in the direction of M81 is ∼ 1.5 mag in the V band.”

1993 April 1 Brian G. Marsden